# ARMOR: High-Performance Semi-Structured Pruning via Adaptive Matrix Factorization

### v1

@@ Abstract
Large language models (LLMs) present significant deployment challenges due to their immense computational and memory requirements. While semi-structured pruning, particularly 2:4 sparsity, offers a path to practical hardware acceleration, existing methods often incur substantial performance degradation. To bridge this gap, we introduce ARMOR: (Adaptive Representation with Matrix-factORization), a novel one-shot post-training pruning algorithm. Instead of directly pruning weights, ARMOR factorizes each weight matrix into a 2:4 sparse core wrapped by two low-overhead, block diagonal matrices. These wrappers act as efficient pre- and post-transformation error correctors, offering greater flexibility to preserve model quality compared to conventional 2:4 pruning techniques. The sparse core and block diagonal wrappers are chosen through a block coordinate descent algorithm that minimizes a layer-wise proxy loss. We theoretically prove this optimization is guaranteed to converge to a solution with a proxy loss less than or equal to state-of-the-art pruning algorithms. Experiments on Llama (Touvron et al., 2023; Dubey et al., 2024) and Qwen (Yang et al., 2025) model families demonstrate that ARMOR consistently and significantly outperforms state-of-the-art 2:4 pruning methods across a wide range of downstream tasks and perplexity evaluations. ARMOR achieves this superior performance while retaining the inference speedups and substantial memory usage reductions of 2:4 pruning, establishing a more effective trade-off between model compression and task accuracy.
@@ 1 Introduction
At the other end of the spectrum, unstructured pruning offers the highest flexibility by removing individual weights from any part of the model. Leading unstructured pruning algorithms have shown that it is possible to remove up to 50% of the weights from an LLM with minimal loss in performance (Frantar & Alistarh, 2023). However, the resulting irregular, sparse matrices disrupt the parallel processing capabilities of modern hardware like GPUs, which are optimized for dense matrix operations. Thus it is difficult to translate the theoretical model size reductions into practical inference speedups is difficult (Xia et al., 2023).
@@
Through extensive experiments on Llama (Dubey et al., 2024; Touvron et al., 2023) and Qwen (Yang et al., 2025) family models, we show that ARMOR consistently and significantly outperforms existing 2:4 pruning methods on both perplexity and downstream tasks. For example, on Llama-2-13B, ARMOR reduces the perplexity gap between the 2:4 pruned model and the dense original by almost 50%. We show that these accuracy gains are achieved while preserving the practical inference speedups and memory reduction inherent to native 2:4 sparsity. Our work suggests that rethinking the fundamental representation of weights, rather than simply removing them, is a promising direction for future hardware-software co-design in efficient deep learning.
@@ 1.1 Related Work
Existing one-shot unstructured/semi-structured pruning methods largely formulate the compressed weight matrix $\hat{W}$ as the element-wise product of a dense matrix $W^{\prime}$ and a binary mask $M$, where $\hat{W}=W^{\prime}\odot M$. Research within this paradigm focuses on two primary challenges: identifying an optimal mask $M$ (importance scoring) and updating the unpruned weights in $W^{\prime}$ to compensate for the removed connections. The simplest approaches are weight-update-free methods like Wanda, which fix the unpruned weights to their original values ($W^{\prime}=W$) and focus on finding a good mask M through element-wise metrics (Sun et al., 2024; Liu et al., 2025; ; Dong et al., 2024; Das et al., 2023; Zhang et al., 2024).
In contrast, more complex weight-update methods such as SparseGPT (Frantar & Alistarh, 2023) aim for higher accuracy by iteratively pruning weights while simultaneously adjusting the remaining ones to minimize a proxy reconstruction loss, often based on a Hessian sketch. While effective, this introduces significant computational overhead, such as the costly inversion of the Hessian sketch matrix. Both methods suffer significantly increased performance loss when applied to 2:4 semi-structured pruning compared with their unstructured counterparts. This has motivated explorations into alternative pruning formulations beyond simple element-wise masking. These include factorization-based methods like DSF (Boža & Macko, 2024), which represents the weight matrix as a product of two sparse matrices, and decomposition-based methods like WRP (Tan et al., 2024), which represents it as a sum of a semi-structured and an unstructured sparse matrix.
However, these alternative approaches face significant practical barriers that prevent theoretical gains from translating into real-world inference efficiency. DSF, for instance, is incompatible with hardware-accelerated 2:4 semi-structured sparsity, as the product of two 2:4 matrices yields no compression over a dense equivalent. Similarly, WRP’s performance is bottlenecked by its reliance on a highly unstructured sparse matrix component, for which no efficient matrix-multiplication kernels exist. Consequently, a gap remains between achieving high compression rates and realizing tangible speedups on modern hardware.
@@ 3.1 ARMOR Matrix Factorization
The key to ARMOR is the diagonal matrix wrappers, $A$ and $B$, that surround the sparse “core” $W^{\prime}\odot M$. Compared to the naive approach of directly pruning, these diagonal matrix wrappers offer additional flexibility compared with the naive sparse-core-only approach, while having low overhead and existing implementations on hardware for storage and inference. We can store $A$ and $B$ as tensors of size $(d_{\mathrm{out}}/d_{\mathrm{block}})\times d_{\mathrm{block}}\times d_{\mathrm{block}}$ and $(d_{\mathrm{in}}/d_{\mathrm{block}})\times d_{\mathrm{block}}\times d_{\mathrm{block}}$ respectively. Matrix multiplication at inference time can be performed as batched matrix multiplication. As a result the overhead of storing $A$ and $B$ and performing inference grows with $\mathcal{O}((d_{\mathrm{out}}+d_{\mathrm{in}})d_{\mathrm{block}})$, which is sublinear to the number of original parameters in the layer, $d_{\mathrm{out}}d_{\mathrm{in}}$.
@@ 3.3 Optimization Algorithm
$(A)_{0}=I,\quad(B)_{0}=I,\quad(W^{\prime})_{0}=\bar{W},\\ (M_{ij})_{0}=\mathbb{I}\left(W_{i,j}\in\text{top}_{2}\left({W}_{i,\left[c\right]}\right)\right)\forall i\in[d_{\mathrm{out}}],j\in[d_{\mathrm{in}}],$ (3)
where $c=\lceil j/4\rceil$ and $I$ denotes the identity matrix of appropriate size. This initialization is chosen as it is the optimal solution to equation 2 if $\hat{W}$ only consisted of the naive sparse core, and is equivalent to the pruning result of NoWag-P pruning algorithm (Liu et al., 2025).
@@ 4 Results
Models and Experimental Setup
We demonstrate the efficacy of our pruning method on two contemporary model families: Qwen 2.5 (7B, 14B, 32B, and 72B) and Qwen 3 (8B and 14B) (Yang et al., 2025). Our investigation targets foundational base models of 7B parameters and larger. This focus on base models allows for a direct assessment of our algorithm’s effect on the core knowledge acquired during pre-training, removing confounding variables introduced by instruction tuning or other post-training modifications. Furthermore, we concentrate on dense architectures, excluding Mixture-of-Experts (MoE) variants, as recent work suggests that MoE models benefit from specialized pruning strategies (Xie et al., 2024; Li et al., 2025b). For all pruning experiments, we configured the block size, $d_{\mathrm{block}}$, to 128 and executed the proxy loss optimization for 20,000 iterations. Additional implementation details are provided in Appendix G.
Evaluation
To comprehensively assess performance degradation, we employed a two-pronged evaluation strategy. First, to measure practical performance on downstream tasks, we evaluated the pruned Qwen models on a suite of seven industry-standard benchmarks using the LM Eval Harness (Gao et al., 2024). These benchmarks cover a range of capabilities, including commonsense and complex reasoning, mathematical problem-solving, and world knowledge. A detailed description of each benchmark is available in Appendix F. Second, to ensure comparability with the broader model compression literature, which often relies on perplexity metrics, we conducted an additional set of experiments. For this, we pruned models from the Llama-2 (7B, 13B, and 70B) (Touvron et al., 2023) and Llama-3 (8B and 70B) (Dubey et al., 2024) families. We then evaluated their perplexity on the test split of Wikitext2 (Merity et al., 2016) and a subset of the C4 validation split (Dodge et al., 2021), following standard evaluation protocols in the field.
Baselines
We compare ARMOR against 3 leading pruning methods, SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), NoWag-P (Liu et al., 2025). Using NoWag-P is of particular interest, since as discussed previously, ARMOR uses the same proxy loss and is initalized at NoWag-P. Therefore comparing ARMOR against NoWag-P servers as an ablation to evaluate the empirical effectiveness of the ARMOR factorization and optimization algorithm.
@@ 4.3 Inference Efficiency
To quantify the real world performance efficiency benefits of ARMOR pruned models, we benchmarked the generation speed, max VRAM, and model size of the original, 2:4 pruned, and ARMOR pruned Qwen 2.5 7B and 14B parameter models. The results are shown in Table 4. Additionally, we benchmarked the individual batched (batch size of 8192) Matrix Vector timing for a standard layer of Qwen 2.5-14B comparing a dense matrix, 2:4 sparse matrix, and ARMOR factorized. Both experiments confirm that the ARMOR block diagonal matrices carry only a minor overhead, enabling ARMOR to retrain the speedup, model size reduction, and max VRAM reduction of properties of 2:4 pruning.
@@ 4.4 Ablations
Figure 3: Left: Relative average Proxy Loss and C4 Perplexity of Llama-2 7B across 20,000 iterations of the ARMOR Proxy Loss optimization algorithm with block size 128. Right: Relative C4 Perplexity for Lama-2 7B/13B, and Llama-3 8B across block sizes of 1, 8, 16, 32, 64, and 128. Each block size was only optimized for 5000 iterations due to time constraints. Relative perplexity is with respect to initial and optimal (dense) perplexities.
@@ 6 Reproducibility Statement
We are committed to ensuring the reproducibility of our results, the source code for this project is currently being prepared for public release and will be made available soon. To facilitate replication, we have also provided detailed descriptions of our model architecture, algorithms, hyperparameter settings, and experimental setup in Appendix G.
@@ Appendix A Notation in Depth
where $D^{\left(i\right)}\in\mathbb{R}^{d_{\mathrm{block}}\times d_{\mathrm{block}}}\quad\forall\quad i\in[1,d/d_{\mathrm{block}}]$. More generally for a matrix $C\in\mathbb{R}^{d_{1}\times d_{2}}$ we denote the $d_{\mathrm{block}}\times d_{\mathrm{block}}$ matrix blocks as:
@@ B.1 The Sparse Core Update Step in Depth
The 2:4 semi-structured sparsity allows for efficient calculation of equations 7 and 8. Since there are only 2 unmasked values for each group $\left(B^{\prime}D^{\left(j\right)}B^{\prime T}\right)$ and $\left(B^{\prime}D^{\left(j\right)}\Delta W^{T}a\right)$ are of shape $2\times 2$ and $2\times 1$. Thus the computational cost for solving for $\left(B^{\prime}D^{\left(j\right)}B^{\prime T}\right)^{\dagger}\left(B^{\prime}D^{\left(j\right)}\Delta W^{T}a\right)$ are negligible. Rather, calculating equations 7 and 8 are dominated by calculating $\left(B^{\prime}D^{\left(j\right)}B^{\prime T}\right)$ and $\left(B^{\prime}D^{\left(j\right)}\Delta W^{T}a\right)$, which scale with $d_{\mathrm{block}}^{2}$. We update all $d_{\mathrm{in}}d_{\mathrm{out}}/d_{\mathrm{block}}^{2}$ blocks at once, thus the overall computational complexity is $\mathcal{O}(d_{\mathrm{in}}d_{\mathrm{out}})$, scaling linearly with parameter size.

### v2

@@ Abstract
Large language models (LLMs) present significant deployment challenges due to their immense computational and memory requirements. While semi-structured pruning, particularly 2:4 sparsity, offers a path to practical hardware acceleration, existing methods often incur substantial performance degradation. To bridge this gap, we introduce ARMOR: (Adaptive Representation with Matrix-factORization), a novel one-shot post-training pruning algorithm. Instead of directly pruning weights, ARMOR factorizes each weight matrix into a 2:4 sparse core wrapped by two low-overhead, block diagonal matrices. These wrappers act as efficient pre- and post-transformation error correctors, offering greater flexibility to preserve model quality compared to conventional 2:4 pruning techniques. The sparse core and block diagonal wrappers are chosen through a block coordinate descent algorithm that minimizes a layer-wise proxy loss. We prove this optimization is guaranteed to converge to a solution with a proxy loss less than or equal to state-of-the-art pruning algorithms. Experiments on Llama (Touvron et al., 2023; Dubey et al., 2024) and Qwen (Yang et al., 2025) model families demonstrate that ARMOR consistently and significantly outperforms state-of-the-art 2:4 pruning methods across a wide range of downstream tasks and perplexity evaluations, and generalizes to provide improvements for general N:M patterns and unstructured sparsity. ARMOR achieves this superior performance while retaining the inference speedups and substantial memory usage reductions of 2:4 pruning, establishing a more effective trade-off between model compression and task accuracy.
@@ 1 Introduction
At the other end of the spectrum, unstructured pruning offers the highest flexibility by removing individual weights from any part of the model. Leading unstructured pruning algorithms have shown that it is possible to remove up to 50% of the weights from an LLM with minimal loss in performance (Frantar & Alistarh, 2023). However, the resulting irregular, sparse matrices disrupt the parallel processing capabilities of modern hardware like GPUs, which are optimized for dense matrix operations. Thus it is difficult to translate the theoretical model size reductions into a practical inference speedup (Xia et al., 2023).
@@
Through extensive experiments on Llama (Dubey et al., 2024; Touvron et al., 2023) and Qwen (Yang et al., 2025) family models, we show that ARMOR consistently and significantly outperforms existing 2:4 pruning methods on both perplexity and downstream tasks. For example, on Llama-2-13B, ARMOR reduces the perplexity gap between the 2:4 pruned model and the dense original by almost 50%. We show that these accuracy gains are achieved while preserving the practical inference speedups and memory reduction inherent to native 2:4 sparsity, and that ARMOR generalizes to general N:M and unstructured sparsity with consisted improvements over SOTA. Our work suggests that rethinking the fundamental representation of weights, rather than simply removing them, is a promising direction for future hardware-software co-design in efficient deep learning.
@@ 1.1 Related Work
Existing one-shot unstructured/semi-structured pruning methods largely formulate the compressed weight matrix $\hat{W}$ as the element-wise product of a dense matrix $W^{\prime}$ and a binary mask $M$, where $\hat{W}=W^{\prime}\odot M$. Research within this paradigm focuses on two primary challenges: identifying an optimal mask $M$ (importance scoring) and updating the unpruned weights in $W^{\prime}$ to compensate for the removed connections. The simplest approaches are weight-update-free methods like Wanda, which fix the unpruned weights to their original values ($W^{\prime}=W$) and focus on finding a good mask M through element-wise metrics (Sun et al., 2024; Liu et al., 2025; ; Dong et al., 2024; Das et al., 2023; Zhang et al., 2024). In contrast, more complex weight-update methods such as SparseGPT (Frantar & Alistarh, 2023) aim for higher accuracy by iteratively pruning weights while simultaneously adjusting the remaining ones to minimize a proxy reconstruction loss, often based on a Hessian sketch. While effective, this introduces significant computational overhead, such as the costly inversion of the Hessian sketch matrix. Both methods suffer significantly increased performance loss when applied to 2:4 semi-structured pruning compared with their unstructured counterparts.
[…]
@@ 3.1 ARMOR Matrix Factorization
The key to ARMOR is the diagonal matrix wrappers, $A$ and $B$, that surround the sparse “core” $W^{\prime}\odot M$. Compared to the naive approach of directly pruning or diagonal only wrappers, these block diagonal matrix wrappers offer additional flexibility, while having low overhead and existing implementations on hardware for storage and inference. We can store $A$ and $B$ as tensors of size $(d_{\mathrm{out}}/d_{\mathrm{block}})\times d_{\mathrm{block}}\times d_{\mathrm{block}}$ and $(d_{\mathrm{in}}/d_{\mathrm{block}})\times d_{\mathrm{block}}\times d_{\mathrm{block}}$ respectively. Matrix multiplication at inference time can be performed as batched matrix multiplication. As a result the overhead of storing $A$ and $B$ and performing inference grows with $\mathcal{O}((d_{\mathrm{out}}+d_{\mathrm{in}})d_{\mathrm{block}})$, which is sublinear to the number of original parameters in the layer, $d_{\mathrm{out}}d_{\mathrm{in}}$.
@@ 3.3 Optimization Algorithm
$(A)_{0}=I,\quad(B)_{0}=I,\quad(W^{\prime})_{0}=\bar{W},\\ (M_{ij})_{0}=\mathbb{I}\left(\mathcal{I}_{i,j}\in\text{top}_{2}\left({\mathcal{I}}_{i,\left[c\right]}\right)\right)\forall i\in[d_{\mathrm{out}}],j\in[d_{\mathrm{in}}],$ (3)
where $c=\lceil j/4\rceil$, $\mathcal{I}_{i,j}=\bar{W}_{i,j}^{2}\|X_{j}\|_{2}^{2}$, and $I$ denotes the identity matrix of appropriate size. This initialization is chosen as it is the optimal solution to equation 2 if $\hat{W}$ only consisted of the naive sparse core, and is equivalent to the pruning result of NoWag-P pruning algorithm (Liu et al., 2025).
@@ 4 Results
Models and Experimental Setup We demonstrate the efficacy of our pruning method on two contemporary model families: Qwen 2.5 (7B, 14B, 32B, and 72B) and Qwen 3 (8B and 14B) (Yang et al., 2025). Our investigation targets foundational base models of 7B parameters and larger. This focus on base models allows for a direct assessment of our algorithm’s effect on the core knowledge acquired during pre-training, removing confounding variables introduced by instruction tuning or other post-training modifications. Furthermore, we concentrate on dense architectures, excluding Mixture-of-Experts (MoE) variants, as recent work suggests that MoE models benefit from specialized pruning strategies (Xie et al., 2024; Li et al., 2025b). For all pruning experiments, we configured the block size, $d_{\mathrm{block}}$, to 128 and executed the proxy loss optimization for 20,000 iterations. Additional implementation details are provided in Appendix H.
Evaluation: To comprehensively assess performance degradation, we employed a two-pronged evaluation strategy. First, to measure practical performance on downstream tasks, we evaluated the pruned Qwen models on a suite of seven industry-standard benchmarks using the LM Eval Harness (Gao et al., 2024). These benchmarks cover a range of capabilities, including commonsense and complex reasoning, mathematical problem-solving, and world knowledge. A detailed description of each benchmark is available in Appendix G. Second, to ensure comparability with the broader model compression literature, which often relies on perplexity metrics, we conducted an additional set of experiments. For this, we pruned models from the Llama-2 (7B, 13B, and 70B) (Touvron et al., 2023) and Llama-3 (8B and 70B) (Dubey et al., 2024) families. We then evaluated their perplexity on the test split of Wikitext2 (Merity et al., 2016) and a subset of the C4 validation split (Dodge et al., 2021), following standard evaluation protocols in the field.
Baselines We compare ARMOR against 3 leading pruning methods, SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), NoWag-P (Liu et al., 2025). Using NoWag-P is of particular interest, since as discussed previously, ARMOR uses the same proxy loss and is initalized at NoWag-P. Therefore comparing ARMOR against NoWag-P servers as an ablation to evaluate the empirical effectiveness of the ARMOR factorization and optimization algorithm.
@@ 4.2 Perplexity Based Evaluations
4.3 Comparisons With Learnable Baselines
We further compare ARMOR against recent learnable semi-structured pruning methods: RotPruner (Chen & Wang, 2025) and DenoiseRotator (Gu et al., 2025). For these specific comparisons, we adopt the baselines’ evaluation protocol: reporting Wikitext-2 perplexity at a sequence length of 2048 (in contrast to the native 4096/8192 used in our main results in Table 3). We exclude MaskLLM (Fang et al., 2024) as it operates in a retraining regime requiring orders of mangitude more data and compute than ARMOR. We also omit Targeted Low-Rank Refinement (Shen et al., 2025) due to the lack of publicly available code. As shown in Table 6, ARMOR significantly outperforms RotPruner across all models. Against the concurrent work DenoiseRotator, ARMOR consistently outperforms the Wanda-based variant and remains highly competitive with the SparseGPT variant, surpassing it on Llama-2 7B and 13B. Crucially, ARMOR achieves this parity without the rigid inference constraints of rotation-based methods. While DenoiseRotator enforces a fixed computational overhead via orthogonal matrices, ARMOR offers a tunable trade-off between accuracy and latency through the block size hyperparameter $d_{\mathrm{block}}$.
@@ 4.4 Inference Efficiency
[…]
4.5 Extensions To N:M and Unstructured
ARMOR extends naturally to general N:M patterns. Furthermore, while the sparse core update is not computationally tractable for general unstructured sparsity, ARMOR can be applied to unstructured sparsity as well by only performing the continuous update step. In table 6 we present experiments comparing ARMOR with NoWag-P for 50% unstructured, 4:8, 5:8, and 6:8 patterns Llama-2 7B/13B. ARMOR consistently and significantly outperforms NoWag-P across all sparsity patterns and types. These experiments were ran with significantly less iterations, 5000 for unstructured, 2000 for semi-structured, than the main 2:4 results, as a result they only serve as a lower bound on ARMOR’s performance on sparsity structures beyond 2:4 sparsity.
Wikitext 2 $(\downarrow)$
Method 2-7B 2-13B 2-70B 3-8B
Dense 5.47 4.88 3.32 6.13
RotPruner 9.20 - - 11.65
Wanda +   DenoiseRotator 9.88 8.77 4.85 11.41
SparseGPT +   DenoiseRotator 8.70 6.81 4.64 10.01
ARMOR 7.65 6.80 4.80 11.27
Table 5: Comparison of Wikitext 2 perplexity (context length = 2048) of 2:4 pruned Llama-2 7B/13B/70B and Llama-3 8B between ARMOR, RotPruner and DenoiseRotator methods. Same ARMOR models as Table 3 evaluated at shorter context length.
Wikitext 2 $(\downarrow)$ C4 $(\downarrow)$
Method Sparsity 2-7B 2-13B 2-7B 2-13B
[…]
NoWag-P 4:8 8.04 6.47 10.17 8.67
ARMOR 4:8+$o$ 6.74 5.96 8.78 7.98
NoWag-P 5:8 5.89 5.14 7.61 6.80
ARMOR 5:8+$o$ 5.62 5.02 7.29 6.67
NoWag-P 6:8 5.35 4.76 6.92 6.27
ARMOR 6:8+$o$ 5.26 4.72 6.80 6.23
Table 6: Comparison of ARMOR vs. NoWag-P across 50% and various N:M sparsity patterns on Llama-2 7B/13B. Relative overhead $o$ is the same as Table 3.
@@ 4.6 Ablations
Figure 3: Left: Relative average Proxy Loss and C4 Perplexity of Llama-2 7B across 20,000 iterations of the ARMOR Proxy Loss optimization algorithm with block size 128. Right: Relative C4 Perplexity for Lama-2 7B/13B, and Llama-3 8B across block sizes of 1 (NoWag-P), 8, 16, 32, 64, and 128. ARMOR was ran for 5000 iterations for each block size. Relative perplexity is with respect to initial and optimal (dense) perplexities.
@@ 6 Reproducibility Statement
Our code for ARMOR is available on Github at https://github.com/LawrenceRLiu/ARMOR/tree/main. To facilitate replication, we have also provided detailed descriptions of our architecture, algorithms, hyperparameter settings, and experimental setup in Appendix H.
@@ Appendix A Notation in Depth
where $D^{\left(i\right)}\in\mathbb{R}^{d_{\mathrm{block}}\times d_{\mathrm{block}}}\quad\forall\quad i\in[1,d/d_{\mathrm{block}}]$. We choose a block diagonal wrappers over diagonal only because diagonal only offers no additional expressivity because it commutes with the element-wise pruning mask, mathematically:
$\mathrm{diag}\left(a\right)(W’\odot M)\mathrm{diag}\left(b\right)=W^{\prime\prime}\odot M,\quad\forall\quad a\in\mathbb{R}^{d_{\mathrm{in}}}\quad b\in\mathbb{R}^{d_{\mathrm{out}}}$ (5)
where $W^{\prime\prime}=\mathrm{diag}\left(a\right)W^{\prime}\mathrm{diag}\left(b\right)$. In other words, Block-Diagonal matrices are a minimal, hardware acceleratable, structure that enable flexibility, which diagonal only matrices do not.
More generally for a matrix $C\in\mathbb{R}^{d_{1}\times d_{2}}$ we denote the $d_{\mathrm{block}}\times d_{\mathrm{block}}$ matrix blocks as:
@@ B.1 The Sparse Core Update Step in Depth
The 2:4 semi-structured sparsity allows for efficient calculation of equations 8 and 9. Since there are only 2 unmasked values for each group $\left(B^{\prime}D^{\left(j\right)}B^{\prime T}\right)$ and $\left(B^{\prime}D^{\left(j\right)}\Delta W^{T}a\right)$ are of shape $2\times 2$ and $2\times 1$. Thus the computational cost for solving for $\left(B^{\prime}D^{\left(j\right)}B^{\prime T}\right)^{\dagger}\left(B^{\prime}D^{\left(j\right)}\Delta W^{T}a\right)$ are negligible. Rather, calculating equations 8 and 9 are dominated by calculating $\left(B^{\prime}D^{\left(j\right)}B^{\prime T}\right)$ and $\left(B^{\prime}D^{\left(j\right)}\Delta W^{T}a\right)$, which scale with $d_{\mathrm{block}}^{2}$. We update all $d_{\mathrm{in}}d_{\mathrm{out}}/d_{\mathrm{block}}^{2}$ blocks at once, thus the overall computational complexity is $\mathcal{O}(d_{\mathrm{in}}d_{\mathrm{out}})$, scaling linearly with parameter size.